### v1

# Simulating jet launching in binary neutron star mergers with no artificial atmosphere limitations

Abstract
We present general relativistic magnetohydrodynamic simulations of binary neutron star (BNS) mergers, exploring the conditions to launch relativistic outflows compatible with gamma-ray burst (GRB) jets. Employing an unprecedentedly low numerical density floor decreasing as the sixth power of radial distance, we obtain for the first time an evolution where such a floor has no effect on any outflow produced during and after merger. In models forming an accreting black hole (BH)  $\simeq\,25$ ms after merger, incipient jets are launched with terminal Lorentz factors and Poynting-flux luminosities compatible with GRBs. The rather high jet velocities reached in a few tens of ms, when compared to previous simulations adopting a much higher and uniform density floor, show the importance of removing any influence of the latter as the jet evolves toward increasingly large time and spatial scales. In absence of collapse, we find the emergence of much heavier and slower polar outflows, incapable of producing a GRB. We thus favor a BH origin for GRB jets.
I Introduction
The coincident detection of gravitational waves (GWs) and electromagnetic (EM) radiation from the binary neutron star (BNS) merger event GW170817 [1, 2, 3, 4, 5, 6] conclusively demonstrated that BNS mergers can power short gamma-ray burst (GRB) jets. However, modeling jet formation in merger simulations to reliably connect theoretical models to observed GRB-related EM signals remains a challenging endeavor.
Previous numerical studies employing general relativistic magnetohydrodynamic (GRMHD) simulations have shown that incipient jets (i.e. collimated, mildly relativistic, and magnetically dominated outflows) can be effectively launched by black hole-disk (BH-disk) systems formed in merger aftermaths, displaying promising terminal Lorentz factors and Poynting-flux luminosities (e.g., [7, 8, 9, 10]). Yet, the formation of an observationally consistent, mature GRB jet is not guaranteed; it critically depends on the jet’s interaction with the surrounding environment and its further evolution over larger spatial and temporal scales. In exploring the transition from promising BH-disk outflows to fully developed, high Lorentz factor jets, the presence of an excessively high atmospheric density floor employed in simulations represents a major limitation. In this work, we overcome this limitation by conducting GRMHD BNS merger simulations with an unprecedentedly low atmospheric density floor, scaling with radial distance as $\propto\!r^{-6}$. This enables a self-consistent evolution where all relevant outflows—from dynamical ejecta to incipient jets—always remain well above the atmospheric threshold.
In scenarios resulting in BH-disk formation, we obtain incipient jets exhibiting properties compatible with GRB phenomenology, and capable of propagating through the baryon-loaded polar outflows expelled prior to collapse by the highly magnetized massive neutron star (NS) remnant (e.g., [11]). [^1] For sufficiently long NS remnant lifetimes, such pre-collapse outflows can have major impact on jet propagation (e.g., [12] and refs. therein), and be potentially responsible for distinctive, relatively unexplored EM signals preceding the main GRB emission (possibly connected to observed GRB precursors [13, 14, 15]). Even with the relatively short evolution considered (up to 100 ms post-merger), the advantages of employing a low-density atmosphere become evident, particularly when comparing to previous works. Our results reinforce the expectation that, as the evolution proceeds on larger scales, a sufficiently low density atmosphere represents a necessary ingredient for a reliable description of emerging jets and their connection to GRB-related signals.
Additionally, we explore a scenario in which no collapse occurs, thereby probing the viability of the ‘magnetar’ GRB model. Consistent with [11] and a number of subsequent investigations [16, 17, 18, 19, 20, 21, 22], our simulations confirm the natural emergence of collimated outflows. However, such outflows appear incompatible with an incipient GRB jet, supporting the prevalent (but still debated) viewpoint that typical short GRBs likely require BH central engines. Notably, employing a much lower and more realistic atmospheric floor has negligible impact for outflows powerd by massive NS remnants. We conclude that removing the limitations associated with the artificial atmosphere in BNS merger simulations further favors a BH origin for GRB jets produced in this context.
II Physical Models and Numerical Setup
We consider an equal-mass BNS system with individual gravitational masses of $1.3624\,M_{\odot}$, corresponding to a chirp mass of $1.186\,M_{\odot}$ (the same as the GW170817 event [23]). The initial data are generated with the LORENE library, with coordinate separation of 45 km and adopting a piecewise-polytropic approximation of the APR4 equation of state (EOS), as implemented in [24]. For the evolution, we consider a hybrid EOS obtained by adding a thermal component to the ‘cold’ piecewise-polytropic part, with the resulting pressure dependence on rest-mass density and specific internal energy given as
${P(\rho,\epsilon)=P_{\rm cold}(\rho)+(\gamma_{\rm th}-1)(\epsilon-\epsilon_{\rm cold}(\rho))\rho\,.}$ (1)
For rest-mass densities above $10^{10}$ g/cm3, we adopt the widely used adiabatic index $\gamma_{\rm th}\!=\!1.8$ (e.g. [24] and refs. therein), while for $\rho\!<\!10^{10}$ g/cm3 we set $\gamma_{\rm th}\!=\!4/3$. The latter choice is motivated by the fact that realistic temperature-dependent EOS have an effective ideal-gas adiabatic index of 4/3 at those low densities [25] (see Supplemental material).
Dipolar magnetic fields are initialized onto the two NSs at two orbits prior to merger, using the vector potential prescription of [16] with fields also extending outside the NS surface. The field strength at the NS poles is set as $B_{\rm pole}\!\simeq\!10^{15}$ G, corresponding to a total magnetic energy of $E_{\rm mag}\simeq 3\times 10^{49}$ erg at merger time. While these pre-merger magnetic fields are unrealistically strong, they still represent a small perturbation. Such field strengths are chosen to achieve the high magnetization levels expected in the post-merger phase, despite the fact that the corresponding amplification mechanisms are not fully resolved. This is a common approach in studies presenting jet-forming BNS merger simulations (e.g. [8, 10]), which are all limited by insufficient resolution.
Our simulations are performed using the GRMHD code Spritz [26, 27], employing the generalised Lorenz gauge to evolve the vector potential [28], PPM reconstruction along with HLLE Riemann solver, RK4 method for time-stepping, and the publicly available RePrimAnd library for primitive variable recovery [29, 30]. Spacetime evolution follows the BSSN formulation as implemented in the MacLachlan module of the Einstein Toolkit [31]. Our Cartesian grid setup includes twelve refinement levels, with the two finest levels tracking each NS during the inspiral phase. At merger, we transition to fixed mesh refinement, and upon collapse to a BH we cover the latter with a cubic shaped extra level with $\simeq\!9.5$ km side. The computational domain extends up to 51,500 km in positive and negative $x,y,z$ directions.
A key novelty compared to previous work concerns the setting of the artificial atmosphere, i.e. the minimum rest-mass density $\rho_{\mathrm{atm}}$ allowed during evolution. We recall that in Spritz, whenever the evolution would lead to $\rho\!<\!\rho_{\mathrm{atm}}$, density is reset to $\rho_{\mathrm{atm}}$, pressure is reset according to the cold part of the EOS, fluid velocity is set to zero, and magnetic fields remain unchanged. Here, we set $\rho_{\mathrm{atm}}\!=\!\rho^{*}\!=\!6.17\times 10^{4}$ g/cm3 up to a radial distance of $r^{*}\!=\!74$ km, while at larger distances we employ the power-law profile $\rho_{\mathrm{atm}}\!=\!\rho^{*}\,(r/r^{*})^{-6}$. With this setting, no atmospheric floor reset occurs at any time above $\approx\!90$ km distance, leaving all outflows, from dynamical ejecta to the incipient jet, completely unaffected. [^2] Atmospheric densities decreasing with distance were already employed, e.g., in [32] There, $\rho_{\mathrm{atm}}\!\propto\!r^{-3}$ down to 0.166 g/cm3 and then the floor remains constant, resulting in $\rho_{\mathrm{atm}}$ higher than ours above $\sim\!700$ km. At 2000 km distance, it is already $\sim\!1000$ times higher (see comparison in Figure 1, top panel).
We present three different simulations. In case A, in order to explore the scenario in which a BH-disk system powers an incipient jet, we manually induce the collapse of the NS remnant at $\approx\!25$ ms after merger, employing the procedure introduced in [33]. In case B, we consider the non-collapsing scenario. In both cases, evolved up to 75 ms after merger, the finest grid spacing is set to $dx\!=\!177$ m ($88.5$ m in the extra level covering the BH in case A). Case C is a lower resolution version of case A, with finest spacing $dx\!=\!200$ m ($100$ m around the BH), evolved up to 100 ms after merger.
Figure 1: Rest-mass density (top) and radial velocity (bottom) along the positive $x$ and $z$ axes at 20 ms after merger, for case A. Rest-mass density floor profiles in the top panel are shown as a black dashed line for our simulations and a blue dotted line for [32] (see text).
III Results
In all cases, the merger produces outflows of unbound material (i.e. dynamical ejecta) in the form of equatorial tidal tails and the more isotropic shock-driven ejecta. The latter, which may affect at later times the propagation and observational signatures of an emerging jet, are characterized by a very fast low-density tail reaching velocities of $\approx\!0.9\,c$ (see Figure 1). Notably, following dynamical ejecta up to their natural front is made possible by the very low and decaying atmospheric floor.
The merger results in a differentially rotating massive NS. At 20 ms post-merger, a helical magnetic tower has formed, extending to over 500 km along the NS spin axis, where magnetic field strength is $\sim\!10^{14}$ G. This tower originates polar outflows with velocities up to $\simeq\!0.2\,c$. Starting from $\simeq\!25$ ms after merger, the collapse to a BH in case A (and C) hampers the further development of such outflows and leads to a very different evolution with respect to case B, in which no collapse occurs.
In cases A and C, the newly formed BH has a mass of $M_{\rm BH}\!\simeq\!2.3\,M_{\odot}$, a dimensionless spin of $\chi\!\equiv\!J/M^{2}\!\simeq\!0.61$, and is surrounded by an accretion disk of mass $M_{\rm disk}\!\simeq\!0.2\,M_{\odot}$. By the end of the simulation, $M_{\rm BH}\!\simeq\!2.4\,M_{\odot}$, $\chi\!\simeq\!0.65$, $M_{\rm disk}\!\simeq\!0.08\,M_{\odot}$, and the accretion rate has stabilized to $\dot{M}\!\simeq\!0.5\,M_{\odot}$/s. Right after collapse, as matter close to the BH is accreted, polar funnels are rapidly escavated along the spin axis. However, magnetic flux within the helical magnetic tower is in good part preserved and, in a few ms, radial velocities in the polar regions and above $\sim\!100$ km distance start to reverse. This leads to a new bipolar highly collimated outflow developing along the BH spin axis, which then matures into an incipient jet around 40 ms after merger.
In Figure 2, we show the magnetic field configuration along with rest-mass density distribution for case A, at about 25 ms and 50 ms after BH formation. At both times, the average magnetic field strength around 500 km distance along the spin axis is $\sim\!10^{14}$ G, while the (mostly toroidal) field close to the BH and in the accretion disk has a strength in the range $10^{15}-10^{16}$ G. The inset of Fig. 2 gives a meridional view of the magnetization level, in terms of the magnetic energy density to rest-mass density ratio $b^{2}/(2\rho)$. For $b^{2}/(2\rho)\!\gg\!1$, this quantity also gives an estimate of the terminal lorentz factor $\Gamma_{\infty}$. Close to the BH, values $\gtrsim 10^{3}$ are achieved, which is consistent with an outflow capable of evolving into a GRB jet. Similar $\Gamma_{\infty}$ are achieved in case C.
Figure 3 (top-panel) shows the time evolution of the total magnetic energy for the three cases. The initial exponential growth associated with the Kelvin-Helmholtz (KH) instability is followed by further amplification sustained by the massive NS differential rotation, via magnetic winding and the magnetorotational instability (MRI) (e.g., [34] and refs. therein). We caution, however, that our resolution is insufficient to fully capture the KH instability as well as the MRI (and the possible $\alpha\Omega$ dynamo driven by it [19]). While in case B magnetic energy keeps growing up to almost $10^{52}$ erg and then starts to slowly decrease, in cases A and C the collapse leads to a sudden drop by a factor of $\approx\!60$, after which a slowly decreasing trend leads to a final value of $\simeq\!6\times 10^{49}$ erg.
Figure 2: 3D rendering of magnetic field-line structure and rest-mass density (red-to-grey colors) for case A, at about $50$ ms (left) and 75 ms (right) after merger. The inset shows a meridional view of the magnetization $b^{2}/(2\rho)$ at the jet base (vertical scale of up to ±$\pm 250$ km).
Figure 3: Evolution of total magnetic energy (top) and Poynting-flux luminosity across the $r\!=\!295$ km spherical surface (bottom), for all cases. The vertical yellow line marks the merger time, whereas vertical black lines mark the collapse time for case A and C, respectively.
Bottom panel of Figure 3 presents the evolution of the Poynting-flux luminosity $L_{\rm EM}$ across the spherical surface $r\!=\!295$ km. Here, the first few peaks within 8 ms after merger relate to the dynamical ejecta crossing the surface. Then, the developing magnetic tower starts to drive a magnetized outflow, gradually increasing $L_{\rm EM}$. For case B, it reaches $\sim\!10^{51}$ erg/s around 25 ms after merger, and this level is maintined until the end. In case A (and C), the collapse causes a temporary decrease until $\approx\!40$ ms after merger, when the contribution of the incipient jet emerges. In the following evolution, $L_{\rm EM}$ remains rather stationary, with a final value of $\simeq 3\times 10^{50}$ erg/s that is of the order of the power expected from the Blandford-Znajek mechanism [35, 36],
${L_{\rm BZ}\sim 5\!\times\!10^{50}\Bigg{(}\frac{\chi}{0.65}\Bigg{)}^{\!2}\Bigg{(}\frac{M_{\rm BH}}{2.4\,M_{\odot}}\Bigg{)}^{\!2}\Bigg{(}\frac{B_{\rm BH}}{5\!\times\!10^{15}\,{\rm G}}\Bigg{)}^{\!2}\,{\rm erg/s}\,.}$ (2)
Focusing now on the incipient jet of case A, we find a rapid acceleration up to $\sim\!1000$ km distance, where a maximum radial velocity of $v^{r}\!\simeq\!0.77\,c$ ($0.52\,c$) is reached on the north (south) side about $40$ ms post-collapse. Evolving further, the location of the maximum $v^{r}$ stalls around the same distance, while its value lowers to $\simeq\!0.6\,c$ ($0.5\,c$). This stalling reflects the fact that the jet is still making its way across the outer part of the pre-collapse polar outflow, as shown in Figure 4. There, we also notice a temporary switch off in jet power along the southern polar region (see also the bending of the helical magnetic field structure in Figure 2). This is due to an earlier instability at the jet-environment interface, triggering a transient baryon loading episode in the funnel.
Figure 4: Meridional view of rest-mass density (top) and radial velocity (bottom) at 75 ms after merger for cases A, B, C, and 100 ms for case C (left to right).
Case C shows a jet evolution similar to case A, with maximum $v^{r}$ reaching $\simeq\!0.69\,c$ before stalling. In terms of north and south differences, it appears more symmetric (Figure 4). Continuing the evolution from 75 to 100 ms after merger, the maximum velocity starts growing again, reaching $\simeq\!0.74\,c$, while the fastest part of the incipient jet advances further before stalling again towards the end. Overall, we observe an outflow piercing through the environment in a series of finite steps, rather than experiencing a continuous advancement and acceleration. After breaking out of the dense material constituting the pre-collapse polar outflows, it will then be able to fully accelerate, potentially reaching the high Lorentz factors required to power a GRB.
In order to compare with previous work, we consider the results of [8], where the merger of (similar but not identical) BNSs, endowed with dipolar magnetic fields and described with piecewise-polytropic EOS, ends up in a BH-disk system launching an incipient jet. In the most promising model (H4M2.8P), covering $\simeq\!35$ ms after collapse, the highest velocity achieved is $\simeq\!0.64\,c$. At a similar time post-collapse, our case A (C) reaches larger values of $\simeq\!0.77\,c$ ($0.69\,c$), despite having factor $\simeq\!1.9$ (2.2) lower resolution (except for the inner $\simeq\!9.5$ km refinement), factor $\sim\!10$ lower magnetic energy, and factor $\sim\!100$ lower Poynting-flux luminosity. Moreover, in the above model H4M2.8P, the collapse occurs only $\sim\!10$ ms after merger, at which time the incipient jet of cases A and C would have propagated (much more easily) across a 60% less massive environment. In a more recent work [10], the authors presented an even more promising model (ALF2-2.70-PU5.5×1016) having resolution, magnetic energy, and collapse time similar to model H4M2.8P. In this case, the maximum jet velocity is $\simeq\!0.83\,c$ (i.e. 8% more than our case A).
While a direct one-to-one comparison is not possible, we note that simulations in [8, 10] employ a uniform atmospheric floor density of order $\sim\!10^{5}$ g/cm3, which is orders of magnitude larger than in the present work (here, with $\rho_{\mathrm{atm}}\!\propto\!r^{-6}$, the floor is already $\sim\!10^{5}$ times lower at 500 km radius). [^3] Notice also that, in [8, 10], velocity remains unchanged when density is reset to the floor value (and not set to zero as in Spritz). This difference could explain how the cases presented here, despite the much longer collapse time and much lower resolution and magnetic energy, can reach velocities almost as high as the most promising model in [10], and higher than any model in [8].
We further note that rest-mass densities within the incipient jet in cases A and C drop below $\!\rho^{*}\!=\!6.17\times 10^{4}$ g/cm3, implying that a uniform density floor with $\rho_{\mathrm{atm}}\!=\!\rho^{*}$ would have affected the evolution. Going to larger time and spatial scales, after the jet breaks out of the surrounding environment and starts to accelerate to high Lorentz factors, the limitations of a high atmospheric floor would become prohibitive. In order to describe jets that are physically consistent and, at the same time, able to evolve into high-$\Gamma$ outflows compatible with GRBs, it is mandatory that their rest-mass densities remain always above the floor. In this work, we have successfully conducted simulations where this condition is strictly maintained at all times.
In case B, the maximum radial velocities never reach values beyond $0.3\,c$ (Figure 4), despite the factor $\sim\!100$ higher magnetic energy and factor $\sim\!3$ higher $L_{\rm EM}$ compared to cases A and C (Figure 3). Here, the fastest part of the outflow remains $2-3$ orders of magnitude more dense (Figure 4) and thus the very low density floor does not make the difference as for the other cases. The final outcome, similar to what found in [11], corresponds to a slow and heavy outflow that has no chance to evolve into a GRB-like jet. Therefore, employing a much lower and more realistic atmospheric floor further favors (in BNS merger context) a GRB progenitor scenario in which the jet is launched by a BH-disk central engine.
IV Conclusions
In this work, we performed BNS merger simulations with the GRMHD code Spritz, considering cases where the remnant massive NS is either long-lived or collapses into a BH $\simeq\!25$ ms after merger. In the latter scenario, an incipient jet emerges, with terminal Lorentz factor and Poynting-flux luminosity that are compatible with a GRB jet. This result, similar to what obtained in previous work (e.g., [7, 8]), is however not sufficient to claim that the system will be able to power a GRB. Further evolution is necessary, knowing that the fate of the incipient jet will strongly depend on the interaction with the surrounding environment.
When exploring the jet propagation on increasingly large scales, densities within the outflow rapidly decrease and the presence of an atmospheric floor (i.e. a minimum density imposed in the simulation) will eventually become a major limitation, unless the floor level is low enough to always remain well below the system’s physical densities. In the jet-forming BNS merger simulations presented here, the latter condition is fulfilled.
While the evolution only covers about $50-75$ ms after collapse, the velocities reached in our simulations ($0.7-0.8\,c$) show the potential advantage of employing a much lower and more realistic density floor, also in comparison to similar models in the literature. In absence of collapse, on the other hand, the very low floor has minor impact on the emerging polar outflows, which are much denser than the incipient jets of BH origin. In this case, velocities remain limited to $\lesssim\!0.3\,c$, further favoring BHs as GRB central engines.
Different caveats affect our simulations, including the low resolution (and consequent unrealistically strong pre-merger magnetic fields) and the use of simple piecewise-polytropic EOS without neutrino transport. Nonetheless, overcoming the limitations associated with the density floor, accomplished here, is independent from the above aspects and represents one necessary step toward a unified description connecting jet-forming BNS mergers with their GRB-related observational signatures.
Acknowledgements.
The authors thank Wolfgang Kastaun, Michail Chabanov, Lorenzo Ennoggi, and Liwei Ji for insightful discussions. The authors gratefully acknowledge the National Science Foundation (NSF) for financial support from grants AST-2009330, AST-2319326, PHY-2110338, PHY-2409706, OAC-2004044, OAC-2411068 as well as the National Aeronautics and Space Administration (NASA) for financial support from TCAN Grants No. 80NSSC18K1488, 80NSSC24K0100, and the Italian Ministry of Foreign Affairs and International Cooperation (MAECI; Grant No. US23GR08). RC, BG, and AP acknowledge further support by the European Union under NextGenerationEU, Mission 4, Component 1, via the PRIN 2022 Project “EMERGE”, Prot. 2022KX2Z3B (CUP C53D23001150006). This research used resources from the Texas Advanced Computing Center’s (TACC) Frontera and Vista supercomputer allocations (awards AST20021 and PHY20010). Additional resources were provided by the BlueSky, Green Prairies, and Lagoon clusters of the Rochester Institute of Technology (RIT) acquired with NSF grants PHY-2018420, PHY-0722703, PHY-1229173, and PHY-1726215.